# Transferable Model-agnostic Vision-Language Model Adaptation for Efficient Weak-to-Strong Generalization

### v2

@@ 6 Conclusion
We present TransMiter, a transferable model-agnostic adapter that enhances vision-language models without fine-tuning. TransMiter utilizes the model’s prediction to capture adaptation knowledge, enabling seamless knowledge transfer from weaker to stronger models, regardless of architecture or size. With its simple structure and forward-only transfer, TransMiter ensures high computational efficiency. Experimental results demonstrate its exceptional transferability across various models, offering a practical solution for scalable and efficient adaptation.
@@ A.2 Datasets
We use 11 visual recognition datasets in our experiments as follows: ImageNet (Deng et al. 2009), EuroSAT (Helber et al. 2018), DTD (Cimpoi et al. 2014), Food101 (Bossard, Guillaumin, and Van Gool 2014), OxfordPets (Parkhi et al. 2012), FGVCAircraft (Maji et al. 2013), Flowers102 (Nilsback and Zisserman 2008), UCF101 (Soomro, Zamir, and Shah 2012), Caltech101 (Fei-Fei, Fergus, and Perona 2004), StanfordCars (Krause et al. 2013), and SUN397 (Xiao et al. 2010). We use all the unlabeled images in each dataset to train and transfer TransMiter. Specifically, for base-to-novel adaptation transfer experiments, we adopt a transductive setting (Menghini, Delworth, and Bach 2023; Li et al. 2024), where images from both seen and unseen classes are accessible during training. Table 7 provides the data statistics.
@@ A.3 Baselines
Fine-tuning In our experiments, we adopt CoOp (Zhou et al. 2022b) and PromptSRC (Khattak et al. 2023) as the methods for soft prompts learning during fine-tuning. Specifically, when transferring CoOp (Zhou et al. 2022b) prompts, we learn a single projection matrix, as learnable prompts are only introduced in the first transformer block of the text encoder. However, since PromptSRC (Khattak et al. 2023) uses deep prompting (khattak et al. 2023), which learns separate prompts at different transformer blocks of the image and text encoder, we employ separate projection matrices for each layer and encoder.
To ensure training conditions similar to those of TransMiter, we adapt the transferred soft prompts to the pre-trained strong VLM in an unsupervised training for one epoch using knowledge distillation, with the fine-tuned weak VLM serving as the teacher model. We set the learning rate to 0.002 and 0.0025 using SGD with a cosine annealing schedule for CoOp (Zhou et al. 2022b) and PromptSRC (Khattak et al. 2023), respectively. For all datasets, we use a batch size of 4, except for Imagenet (Deng et al. 2009) where we used a batch size of 16.
For the other supervised fine-tuning methods compared and applied in Table 3, such as Maple (khattak et al. 2023), HPT (Wang et al. 2024), and LwEIB (Yang et al. 2025), we use ViT-L/14 as the base model and strictly follow the original hyperparameter configurations described in each paper and codebase.
@@ A.4 Implementation details of TransMiter
The MLP layer in TransMiter consists of [layernorm-linear-activation-linear], with zero-initialization applied to the last linear layer to preserve the pre-trained knowledge. We employ orthogonal parameterization techniques (Miles, Elezi, and Deng 2024) using matrix exponentiation, removing the need for additional orthogonality constraints in the loss function. For the hyperparameters, we set the dimensionality of the latent space $D$ and logits $M$ to 1024, with the bottleneck dimensionality of the MLP to four times that of the latent space as default. To train TransMiter, we set the learning rate and weight decay to $1\times 10^{-3}$ and trained for 10 epochs on all datasets, using the AdamW (Loshchilov and Hutter 2019) optimizer. Since TransMiter does not require access to the internal VLM parameters, we extract image and text features before training to reduce computational costs. For the data augmentation, we applied Gaussian noise to the input logits. The temperature $\tau_{\text{ut-s}}$ and $\tau_{\text{ft-s}}$ were set to 0.01 and 0.005, respectively. The regularization weight $\beta$ is set to 500 as default.
@@ Appendix B Additional experiments
In this section, we conduct additional experiments including (1) additional baseline and model combination, (2) Transfer across different pre-training schemes, (3) extreme low shot settings, (4) domain generalization, (5) application on the larger model and (6) application on object detection . We utilize ViT-B/16 and ViT-L/14 as the source and target models, respectively, with PromptSRC (Khattak et al. 2023) as the default fine-tuning strategy unless specified.
@@ B.1 Additional baseline and model combination
In this section, we conduct detailed ablations and analyses of TransMiter, including (1) ablation on a basis change, (2) ablation on a transition matrix, (3) auxiliary class sampling strategy, (4) latent dimensionality, (5) regularization weight (6) analysis on adaptation knowledge, (7) component-wise transfer time analysis, and (8) comparison with PromptKD.
C.1 Ablation on basis change.
@@
C.2 Ablation on a transition matrix.
@@
Sampling Base Novel HM
Random 85.27 81.07 83.12
FPS 85.39 81.14 83.21
Top-k 84.98 80.98 82.93
Table 16: Auxiliary class sampling strategy.
$D$ 512 1024 2048
HM 82.97 83.12 82.69
Table 17: Latent dimensionality $D$.
C.3 Ablation on auxiliary class sampling strategy.
C.4 Sensitivity analysis on latent dimensionality.
@@
C.5 Sensitivity analysis on Regularization weight.
@@
C.6 Analysis on adaptation knowledge.
@@
C.7 Component-wise transfer time analysis
@@
C.8 Comparison with PromptKD.
@@
TransMiter w/o BS 0 0 285 77.05
TransMiter w/ BS 83 3784 285 77.11
Table 20: Efficiency analysis on the ImageNet. BS refers to Basis Change.
@@
We also compared TransMiter with the other state-of-the-art unsupervised adaptation method, PromptKD (Li et al. 2024) in Table 19 and Table 20. PromptKD is similar to Prompt Transfer in that it transfers knowledge from a fine-tuned teacher model to a student model using naive knowledge distillation (Hinton, Vinyals, and Dean 2015). Additionally, it involves transferring specific parameters of the teacher model, such as the text encoder, which were obtained during the fine-tuning. We strictly adhere to the implementation details provided in PromptKD, with the only difference being that we use a weaker model (e.g., ViT-B/16) as the teacher and a stronger model (e.g., ViT-L/14) as the student.
As shown in Table 19, TransMiter significantly outperforms PromptKD across all metrics and adaptation transfer settings. Notably, PromptKD incurs significant transfer time and memory costs due to the need for backpropagation through internal VLM parameters, whereas TransMiter requires negligible or even zero transfer costs. Additionally, TransMiter enables efficient inference, consisting of only a few stacked layers, whereas deep prompting (khattak et al. 2023) slows inference in PromptKD. These results further emphasize that, within the “knowledge extraction and transfer” framework, TransMiter facilitates knowledge transfer from weaker to stronger models more effectively and efficiently than PromptKD.

### v3

@@ 6 Conclusion
We present TransMiter, a transferable model-agnostic adapter that enhances vision-language models without backpropagation. TransMiter utilizes the model’s prediction to capture adaptation knowledge, enabling seamless knowledge transfer from weaker to stronger models, regardless of architecture or size. With its simple structure and forward-only transfer, TransMiter ensures high computational efficiency. Experimental results demonstrate its exceptional transferability across various models, offering a practical solution for scalable and efficient adaptation.
Acknowledgments
This work was partly supported by Korea Research Institute for defense Technology planning and advancement - Grant funded by Defense Acquisition Program Administration(DAPA)(KRIT-CT-23-021) (30%), the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (NRF-2023R1A2C2005373) (25%), Institute for Information & communications Technology Promotion (IITP) grant funded by the Korea government (MSIP) (No. RS-2024-00443251, Accurate and Safe Multimodal, Multilingual Personalized AI Tutors) (25%), and the Korea Health Technology R&D Project through the Korea Health Industry Development Institute (KHIDI), funded by the Ministry of Health & Welfare, Republic of Korea (grant number: HR20C0021) (20%).
@@ A.2 Datasets
We evaluate our method on 11 visual recognition datasets for both base-to-base and base-to-novel adaptation transfer settings, including ImageNet (Deng et al. 2009), EuroSAT (Helber et al. 2018), DTD (Cimpoi et al. 2014), Food101 (Bossard et al. 2014), OxfordPets (Parkhi et al. 2012), FGVCAircraft (Maji et al. 2013), Flowers102 (Nilsback and Zisserman 2008), UCF101 (Soomro et al. 2012), Caltech101 (Fei-Fei et al. 2004), StanfordCars (Krause et al. 2013), and SUN397 (Xiao et al. 2010). For fine-tuning the models, we follow CoOp (Zhou et al. 2022b) and CoCoOp (Zhou et al. 2022a), using a 16-shot setup. All unlabeled image samples (Li et al. 2024) are used for training and transfer in TransMiter. We use all the unlabeled images in each dataset to train and transfer TransMiter. Specifically, for base-to-novel adaptation transfer experiments, we adopt a transductive setting (Menghini et al. 2023; Li et al. 2024), where images from both seen and unseen classes are accessible during training. Table 6 provides the data statistics.
@@ A.3 Baselines
Fine-tuning In our experiments, we adopt CoOp (Zhou et al. 2022b) and PromptSRC (Khattak et al. 2023b) as the methods for soft prompts learning during fine-tuning. Specifically, when transferring CoOp (Zhou et al. 2022b) prompts, we learn a single projection matrix, as learnable prompts are only introduced in the first transformer block of the text encoder. However, since PromptSRC (Khattak et al. 2023b) uses deep prompting (Khattak et al. 2023a), which learns separate prompts at different transformer blocks of the image and text encoder, we employ separate projection matrices for each layer and encoder.
To ensure training conditions similar to those of TransMiter, we adapt the transferred soft prompts to the pre-trained strong VLM in an unsupervised training for one epoch using knowledge distillation, with the fine-tuned weak VLM serving as the teacher model. We set the learning rate to 0.002 and 0.0025 using SGD with a cosine annealing schedule for CoOp (Zhou et al. 2022b) and PromptSRC (Khattak et al. 2023b), respectively. For all datasets, we use a batch size of 4, except for Imagenet (Deng et al. 2009) where we used a batch size of 16.
For the other supervised fine-tuning methods compared and applied in Table 3, such as Maple (Khattak et al. 2023a), HPT (Wang et al. 2024), and LwEIB (Yang et al. 2025b), we use ViT-L/14 as the base model and strictly follow the original hyperparameter configurations described in each paper and codebase.
@@ A.4 Implementation details of TransMiter
The MLP layer in TransMiter consists of [layernorm-linear-activation-linear], with zero-initialization applied to the last linear layer to preserve the pre-trained knowledge. We employ orthogonal parameterization techniques (Miles et al. 2024) using matrix exponentiation, removing the need for additional orthogonality constraints in the loss function. For the hyperparameters, we set the dimensionality of the latent space $D$ and logits $M$ to 1024, with the bottleneck dimensionality of the MLP to four times that of the latent space as default. To train TransMiter, we set the learning rate and weight decay to $1\times 10^{-3}$ and trained for 10 epochs on all datasets, using the AdamW (Loshchilov and Hutter 2019) optimizer. Since TransMiter does not require access to the internal VLM parameters, we extract image and text features before training to reduce computational costs. For the data augmentation, we applied Gaussian noise to the input logits. The temperature $\tau_{\text{ut-s}}$ and $\tau_{\text{ft-s}}$ were set to 0.01 and 0.005, respectively. The regularization weight $\beta$ is set to 500 as default.
@@ Appendix B Additional experiments
In this section, we conduct additional experiments including (1) additional baseline and model combination, (2) Transfer across different pre-training schemes, (3) extreme low shot settings, (4) domain generalization, (5) larger model and (6) object detection . We utilize ViT-B/16 and ViT-L/14 as the source and target models, respectively, with PromptSRC (Khattak et al. 2023b) as the default fine-tuning strategy unless specified.
@@ B.1 Additional baseline and model combination
In this section, we conduct detailed ablations and analyses of TransMiter, including (1) ablation on anchor quantity (2) auxiliary class sampling strategy, (3) latent dimensionality, (4) basis change, (5) transition matrix, (6) regularization weight, (7) analysis on adaptation knowledge, (8) component-wise transfer time, and (9) comparison with PromptKD.
C.1 Effect of anchor quantity.
Model $M$ DTD ($\Delta$) UCF ($\Delta$) Cars ($\Delta$) SUN ($\Delta$)
47 100 196 397
Source $N_{\text{task}}$ 69.82, 0.00 83.84, 0.00 79.62, 0.00 76.28, 0.00
256 71.77, +1.95 84.30, +0.46 79.96, +0.34 -
512 72.32, +2.50 84.32, +0.48 79.87, +0.25 76.07, -0.21
1024 72.54, +2.72 84.37, +0.53 80.52, +0.90 76.46, +0.18
Target $N_{\text{task}}$ 64.18, 0.00 81.96, 0.00 81.38, 0.00 76.52, 0.00
256 71.91, +7.73 84.25, +2.29 81.47, +0.09 -
512 73.64, +9.46 85.15, +3.19 83.03, +1.65 76.79, +0.27
1024 74.21, +10.03 85.52, +3.56 83.11, +1.73 77.14, +0.62
In Table 13, we examine the effects of anchor quantity $M$ on both source and target models within a base-to-base setting. We observe that increasing $M$ enhances the performance of both the source and target models. Notably, the performance gain from additional anchors for the target model is relatively higher across all reported datasets than for the source model. Further, the performance increases with more anchors. These results underscore two main implications: (1) incorporating auxiliary classes enhances representational capacity, leading to more effective training of TransMiter and (2) they play a crucial role in achieving high transferability.
Sampling Base Novel HM
Random 85.27 81.07 83.12
FPS 85.39 81.14 83.21
Top-k 84.98 80.98 82.93
Table 14: Auxiliary class sampling strategy.
$D$ 512 1024 2048
HM 82.97 83.12 82.69
Table 15: Latent dimensionality $D$.
C.2 Ablation on auxiliary class sampling strategy.
C.3 Sensitivity analysis on latent dimensionality.
C.4 Ablation on basis change.
@@
C.5 Ablation on a transition matrix.
@@
C.6 Sensitivity analysis on Regularization weight.
@@
C.7 Analysis on adaptation knowledge.
@@
C.8 Component-wise transfer time analysis
@@
C.9 Comparison with PromptKD.
@@
TransMiter w/o BC 0 0 285 77.05
TransMiter w/ BC 83 3784 285 77.11
Table 20: Efficiency analysis on the ImageNet. BC refers to Basis Change.
@@
We also compared TransMiter with the other state-of-the-art unsupervised adaptation method, PromptKD (Li et al. 2024) in Table 19 and Table 20. PromptKD is similar to Prompt Transfer in that it transfers knowledge from a fine-tuned teacher model to a student model using naive knowledge distillation (Hinton et al. 2015). Additionally, it involves transferring specific parameters of the teacher model, such as the text encoder, which were obtained during the fine-tuning. We strictly adhere to the implementation details provided in PromptKD, with the only difference being that we use a weaker model (e.g., ViT-B/16) as the teacher and a stronger model (e.g., ViT-L/14) as the student.
As shown in Table 19, TransMiter significantly outperforms PromptKD across all metrics and adaptation transfer settings. Notably, PromptKD incurs significant transfer time and memory costs due to the need for backpropagation through internal VLM parameters, whereas TransMiter requires negligible or even zero transfer costs. Additionally, TransMiter enables efficient inference, consisting of only a few stacked layers, whereas deep prompting (Khattak et al. 2023a) slows inference in PromptKD. These results further emphasize that, within the “knowledge extraction and transfer” framework, TransMiter facilitates knowledge transfer from weaker to stronger models more effectively and efficiently than PromptKD.